Tina has saved $20 000 from her summer jobs. Rather than work for a living, she plans to buy an annuity from a trust company and become a beachcomber in Fiji. An annuity will pay her a certain amount each month for the rest of her life and is calculated at 7 percent interest, compounded monthly, over Tina’s 55 remaining years. Tina calculates that she needs at least $5 per day to live in Fiji, and she needs $1200 for air fare. Can she retire now? How much would she have available to spend each day?
Answer #1

Minimum amount Tina will require to live for another 55 years = 5 × 365 × 55 + 1200 = 100375 + 1200 = $ 101575

Number of months in 55 years = 55 × 12 = 660

Total amount with interest rate over 55 years =

20000 (1+0.07)660 = 20000 × (1.07)660

Amount available per day = 20000(1.07)660/ (55×365) = 0.996(1.07)660